# Leveraging Large Language Models for NLG Evaluation: A Survey

### v1

@@ Abstract
In the rapidly evolving domain of Natural Language Generation (NLG) evaluation, introducing Large Language Models (LLMs) has opened new avenues for assessing generated content quality, e.g., coherence, creativity, and context relevance. This survey aims to provide a thorough overview of leveraging LLMs for NLG evaluation, a burgeoning area that lacks a systematic analysis. We propose a coherent taxonomy for organizing existing LLM-based evaluation metrics, offering a structured framework to understand and compare these methods. Our detailed exploration includes critically assessing various LLM-based methodologies, as well as comparing their strengths and limitations in evaluating NLG outputs. By discussing unresolved challenges, including bias, robustness, domain-specificity, and unified evaluation, this survey seeks to offer insights to researchers and advocate for fairer and more advanced NLG evaluation techniques.
@@ 1 Introduction
Organization of this survey: We present the first comprehensive survey of recent advancements in leveraging LLMs for NLG evaluation. Initially, we establish a formal framework for NLG evaluation and propose a taxonomy to categorize relevant works (Section 2). Subsequently, we delve into and elaborate on these works in detail (Section 3). Furthermore, we conduct a systematic review of various meta-evaluation benchmarks that assess the efficacy of LLM-based evaluators (Section 4). In recognition of the rapid evolution of this field, we identify and discuss several potential open problems that may guide future research (Section 5). Concluding this systematic survey, we advocate for the advancement of this field through the development of more impartial, robust, expert and unified LLM-based evaluators. Additionally, we emphasize integrating other evaluation approaches, such as human judgment, to achieve a more comprehensive and multifaceted assessment framework.
@@ 2 Formalization and Taxonomy
SummaC Laban et al. (2022), RiSum Skopek et al. (2023), OpinSummEval Shen and Wan (2023) , leaf, text width=33.6em ] ] [ Dialogue Generation [ FED Mehri and Eskenazi (2020a), Topical-Chat Mehri and Eskenazi (2020b), PersonaChat Mehri and Eskenazi (2020b) , leaf, text width=33.6em ] ] [ Image Caption [ Flickr8K-Expert Hodosh et al. (2013), Composite Aditya et al. (2015), Pascal-50S Vedantam et al. (2015),
@@
In this survey, we classify works in NLG evaluation along three primary dimensions: evaluation task, evaluation references and evaluation function. These dimensions provide a comprehensive framework for categorizing and understanding different approaches within this domain.
@@
Evaluation Task $\mathcal{T}$: NLG encompasses a diverse range of tasks, such as Machine Translation (MT) Farhad et al. (2021); Bapna et al. (2019), Text Summarization (TS) Liu and Liu (2021); Zhang et al. (2023a), Dialogue Generation (DG) Wang et al. (2022); Kann et al. (2022), Story Generation (SG) Yang et al. (2022); Fan et al. (2018), Image Caption generation (IC) Tewel et al. (2022); Zhou et al. (2022), Data-to-Text generation (D2T) Lin et al. (2023); Jing et al. (2023) and General Generation (GE) Wang et al. (2023g); Zheng et al. (2023), each with its unique evaluation requirements and challenges. The specific nature of each task determines the target evaluation aspects and scenarios. For instance, in text summarization, the focus may be on relevance to the source content, while in dialogue generation, the coherence of the response is crucial. Given these varied objectives, our taxonomy also extends to the lens of task-specific evaluation. This categorization enables us to understand how different evaluation methods perform across a spectrum of NLG tasks, thus offering insights into the strengths and limitations of existing evaluation paradigms in distinct task contexts.
@@
Evaluation References $r$: According to whether the references are available, we divide the evaluation scenarios into reference-based and reference-free scenarios. In reference-based evaluation, the generated text $h$ is compared against a set of ground truth references $r$. This approach is particularly prevalent in tasks where the quality of the generated text can be objectively measured against established standards. The comparison metrics might focus on aspects like accuracy, relevance, coherence, and the degree of similarity to the references. Typical applications include text summarization, where the generated summaries are evaluated against reference summaries, and machine translation, where the translations are compared with standard translations. The reference-free approach, in contrast, does not rely on any external references for evaluation. This method evaluates the generated text $h$ based on the intrinsic qualities or its alignment with the provided source context $s$. Evaluation in this context may focus on aspects such as fluency, originality, relevance to the context, etc.
@@
Evaluation Function $f$: Evaluation function could be matching-based or generative-based on the basis of different ways of utilizing LLMs. As shown in Figure 2, matching-based methods measure the semantic equivalence between the reference and hypothesis or measure the proper degree between the source text and hypothesis. Several works measure the semantic equivalence between the reference and hypothesis by using token-level matching functions in distributed representation space Zhang et al. (2020); Zhao et al. (2019) or discrete string space Lin (2004); Papineni et al. (2002). Others focus on sequence-level, such as Sellam et al. (2020); Rei et al. (2020); Peyrard et al. (2017). In contrast, generative-based methods include methods where LLMs are employed to generate evaluation metrics directly. These methods leverage the generative capabilities of LLMs to assess the quality of generated text by designing instructions.
@@
Scope of this survey: Most recently, matching-based methods are based on a neural encoder to calculate a score-specific aspect of evaluation. However, these methods often face challenges such as limited interpretability, lower correlation with human judgments, and a restricted range of evaluated aspects Xu et al. (2023); Fu et al. (2023). Fortunately, the emerging capabilities of LLMs open up a wealth of possibilities for NLG evaluation. This includes improved interpretability through CoT, higher customization via instruction-following capabilities, and better alignment with human evaluations through RLHF Xu et al. (2023); Zheng et al. (2023). Given the abundance of recent surveys primarily focusing on matching-based evaluation methods (refer to Celikyilmaz et al. (2020); Sai et al. (2022); Goyal et al. (2023) for comprehensive summaries), this survey is dedicated to exploring more burgeoning generative-based methods. Figure 3 presents our taxonomy of generative-based evaluation. We classify relevant works into two main categories: prompt-based and tuning-based evaluation, depending on whether the LLM is tuned. Further, we divide these methods into subcategories: score-based, probability-based, likert-style, pairwise comparison, ensemble, and advanced evaluation protocols, each distinguished by their evaluation form. These categories will be detailed in Section 3.
@@ 3 Generative Evaluation
This section delves into a detailed exploration of these two overarching categories of evaluation methods, each accompanied by their respective evaluation protocols. Table 2 provides a comprehensive overview of current prompt-based and tuning-based evaluation methods. The table delineates their respective adaptation tasks, backbone models, scoring protocols, and evaluated aspects for clarity and reference.
@@ 3.1 Prompt-based Evaluation
Prompt-based text evaluation stands at the forefront of advancements in NLG, particularly leveraging the capabilities of LLMs. In this method, the evaluation process is intricately woven into the crafting of prompts – specialized cues designed to guide LLMs in assessing the quality and coherence of generated text. More recently, the Eval4NLP workshop held a shared task on prompting large language models as explainable metrics Leiter et al. (2023). Typically, a prompt template serves as a structured framework that encompasses instructions, aspects, criteria, and desired output formats, providing a systematic guide for evaluating generated text. These templates empower researchers and practitioners to articulate precise evaluation requirements, ensuring consistency and reproducibility in the assessment process. By harnessing the prowess of LLMs, prompt-based evaluation not only provides a comprehensive understanding of NLG system performance but also offers a nuanced approach to extracting valuable insights.
@@
As the actual evaluation process often involves collaborative evaluation by multiple human annotators, some works utilize multiple LLM evaluators with different base models or prompts and allow them to evaluate the quality of generated text from different perspectives, as shown in Figure 5. Wu et al. (2023a) set multiple roles for the LLM to evaluate the quality of the generated summary by comparing it with the reference one on both subjective and objective dimensions. This work generates dynamic role profiles according to input texts and synthesizes the results of multiple roles as the final evaluation result. Li et al. (2023c) utilized multiple LLM evaluators to conduct a pairwise evaluation for the model-generated responses by performing multiple rounds of discussions on the comparison results to reach a mutual agreement on the pairwise scoring. Similarly, Zhang et al. (2023b) proposed to set up LLM evaluators as multiple-layer neural network structures. The bottom evaluators obtain the evaluation result of model-generated responses from a specific evaluation perspective. The upper evaluators receive all evaluation information from the previous layer and discuss it with each other to obtain a more comprehensive evaluation result. Besides, Chan et al. (2023) designed diverse communication strategies with various role prompts during collaborative discussions to evaluate the pairwise model-generated responses.
@@
Some recent works investigate advanced evaluation techniques aimed at achieving more thorough and nuanced assessment outcomes by leveraging chain-of-thought, in-context learning capabilities, fine-grained analysis, etc. A representative fine-grained evaluation method is shown in Figure  4. Liu et al. (2023c) utilized LLMs with chain-of-thought (CoT) prompting and a form-filling paradigm to evaluate the quality of generated texts across various NLG tasks. Min et al. (2023) proposed a find-grained evaluation schema that first extracts a series of atomic facts from the LLM-generated long text, and then utilizes LLM evaluator to validate each atomic fact with the given knowledge source. Lu et al. (2023) proposed a new prompting method called Error Analysis Prompting (EAPrompt) that combines CoT to prompt the LLM evaluator to analyze different types of pre-defined errors (e.g., major and minor errors) in the generated translation based on the given source text and reference translation, and then measures the quality of a generated translation with the previous error analysis. To enhance and improve the robustness of LLM-based evaluators, Hasanbeig et al. (2023) proposed ALLURE, a systematic protocol for auditing and improving LLM-based evaluation of text using iterative in-context-learning. Considering that the evaluation with a single or few references may not accurately reflect the quality of the model’s hypotheses, Tang et al. (2023) leveraged large language models (LLMs) to paraphrase a single reference into multiple high-quality ones in diverse expressions, which enhances various evaluation methods on machine translation, text summarization, and image caption tasks. To further task advantages of the in-context learning capability of LLMs, Liu et al. (2023f) proposed AUTOCALIBRATE to automatically align and calibrate an LLM-based evaluator through incorporating the mined and calibrated rubrics into scoring instructions.
@@ 3.2 Tuning-based Evaluation
In the ever-evolving landscape of Natural Language Generation (NLG) evaluation, a noteworthy paradigm shift is underway as researchers increasingly turn their attention towards fine-tuning open-source language models (e.g., LLaMA), in lieu of traditional closed-based LLMs like ChatGPT and GPT-4. This transformative shift is propelled by a thorough exploration of key perspectives, including the expenses associated with API calls, the robustness of prompting, and the pivotal consideration of domain adaptability.
@@
Compared with prompting-based evaluation, tuning-based methods can also be categorized into various types based on their scoring protocol, such as likert-style evaluation, probability-based evaluation and pairwise evaluation. In addition, based on the output explanations in supervised fine-tuning, these methods can be further divided into holistic evaluation or error-oriented evaluation. We will begin by introducing various types of scoring protocols and subsequently provide a summary of two output explanations in the final advance evaluation.
@@
Some works tune LLMs to provide quality ratings or labels for generated texts. Gekhman et al. (2023) employed FLAN-PaLM 540B Chung et al. (2022) to annotate the quality of real model-generated summaries and utilized these annotated data as training data to tune a light-weight LLM (e.g., T5-11B) as a factual consistency summary evaluator, which predicts “1" if the summary demonstrates factual consistency and “0" otherwise. Yue et al. (2023) reused and repurposed the existing fact-checking, NLI, and summarization tasks datasets and obtained simulated data from open-domain QA datasets to tune light-weight LLMs for attribution evaluation, which generates attributable, extrapolatory or contradictory labels for the generated answer with given query and reference documents. Li et al. (2023a) created a dataset containing multiple scenarios and used GPT-4 OpenAI (2023) to generate evaluation judgments for each scenario as supervision signals to tune LLaMA as a generative evaluator, which can output overall quality rating for individual LLM-generated response in various scenarios. Wang et al. (2023a) repurposed existing datasets with proper anonymization and new personalized labels to tune LLaMA2 Touvron et al. (2023) as a personalized story evaluation model which provides personalized evaluation for generated texts through outputting a grade in [1, 10] and detailed reviews. Kim et al. (2023a) prompted GPT-4 to construct training data, including reference answers and crafted diverse customized score rubrics, and used them to tune LLaMA to evaluate model-generated responses of given instruction, which is generalized to realistic user demands. Ke et al. (2023) instructed GPT-4 to collect referenced and reference-free training data with dialogue-based prompting, utilized to tune instruction-tuned LLMs for evaluating the alignment of model-generated texts with human instructions through generating scores and explanations. Liu et al. (2023a) constructed a reference-free instruction-tuning dataset tailored for multi-aspect evaluation across summarization, dialogue and data-to-text tasks. Considering that there is an internal correlation between the evaluation aspects, this work tuned with auxiliary aspects additionally on diverse evaluation task forms. During inference, this work combined auxiliary and target aspects and predicted either the “Yes” or “No” label to judge whether the generated text satisfied the target aspect and compute the evaluation score.
@@
Some works train generative LLMs to calculate the generation probability of generated texts to evaluate text quality. Thompson and Post (2020) trained a transformer as a multilingual reference-to-candidate paraphraser to obtain the generated probability of model-generated translation based on their reference texts. Qin et al. (2022) tuned the T5 model in the generative and discriminative fashion, and used the probability of generating a text as the quality score.

### v2

# Leveraging Large Language Models for NLG Evaluation: Advances and Challenges

@@ Abstract
In the rapidly evolving domain of Natural Language Generation (NLG) evaluation, introducing Large Language Models (LLMs) has opened new avenues for assessing generated content quality, e.g., coherence, creativity, and context relevance. This paper aims to provide a thorough overview of leveraging LLMs for NLG evaluation, a burgeoning area that lacks a systematic analysis. We propose a coherent taxonomy for organizing existing LLM-based evaluation metrics, offering a structured framework to understand and compare these methods. Our detailed exploration includes critically assessing various LLM-based methodologies, as well as comparing their strengths and limitations in evaluating NLG outputs. By discussing unresolved challenges, including bias, robustness, domain-specificity, and unified evaluation, this paper seeks to offer insights to researchers and advocate for fairer and more advanced NLG evaluation techniques.
@@ 1 Introduction
Organization of this paper: We present the first comprehensive survey of recent advancements in leveraging LLMs for NLG evaluation. Initially, we establish a formal framework for NLG evaluation and propose a taxonomy to categorize relevant works (Section 2). Subsequently, we delve into and elaborate on these works in detail (Section 3). Furthermore, we conduct a systematic review of various meta-evaluation benchmarks that assess the efficacy of LLM-based evaluators (Section 4). Additionally, we provide a thorough comparison of LLM-based evaluators with traditional evaluators in terms of performance, efficiency and qualitative qualitative analysis (Section 5). In recognition of the rapid evolution of this field, we identify and discuss several potential open problems that may guide future research (Section 6). To conclude, we advocate for the advancement of this field through the development of more impartial, robust, expert and unified LLM-based evaluators.
@@ 2 Formalization and Taxonomy
SummaC Laban et al. (2022), RiSum Skopek et al. (2023), OpinSummEval Shen and Wan (2023) , leaf, text width=33.6em ] ] [ Dialogue Generation [ FED Mehri and Eskenazi (2020a), Topical-Chat Gopalakrishnan et al. (2019), PersonaChat Zhang et al. (2018) , leaf, text width=33.6em ] ] [ Image Caption [ Flickr8K-Expert Hodosh et al. (2013), Composite Aditya et al. (2015), Pascal-50S Vedantam et al. (2015),
@@
In this paper, we classify works in NLG evaluation along three primary dimensions: evaluation task, evaluation references and evaluation function. These dimensions provide a comprehensive framework for categorizing and understanding different approaches within this domain.
Evaluation Task $\mathcal{T}$
@@
: NLG encompasses a diverse range of tasks, such as Machine Translation (MT) Farhad et al. (2021); Bapna et al. (2019), Text Summarization (TS) Liu and Liu (2021); Zhang et al. (2023a), Dialogue Generation (DG) Tao et al. (2018); Kann et al. (2022), Story Generation (SG) Yang et al. (2022); Fan et al. (2018), Image Caption (IC) Tewel et al. (2022); Zhou et al. (2022), Data-to-Text generation (D2T) Lin et al. (2023); Jing et al. (2023) and General Generation (GE) Wang et al. (2023g); Zheng et al. (2023), each with its unique evaluation requirements and challenges. The specific nature of each task determines the target evaluation aspects and scenarios. For instance, in text summarization, the focus may be on relevance to the source content, while in dialogue generation, the coherence of the response is crucial. Given these varied objectives, our taxonomy also extends to the lens of task-specific evaluation. This categorization enables us to understand how different evaluation methods perform across a spectrum of NLG tasks, thus offering insights into the strengths and limitations of existing evaluation paradigms in distinct task contexts.
Evaluation References $r$
@@
: According to whether the references are available, we divide the evaluation scenarios into reference-based and reference-free scenarios. In reference-based evaluation, the generated text $h$ is compared against a set of ground truth references $r$. This approach is particularly prevalent in tasks where the quality of the generated text can be objectively measured against established standards. The comparison metrics might focus on aspects like accuracy, relevance, coherence, and the degree of similarity to the references. Typical applications include text summarization, where the generated summaries are evaluated against reference summaries, and machine translation, where the translations are compared with standard translations. The reference-free approach, in contrast, does not rely on any external references for evaluation. This method evaluates the generated text $h$ based on the intrinsic qualities or its alignment with the provided source context $s$. Evaluation in this context may focus on aspects such as fluency, originality, relevance to the context, etc.
Evaluation Function $f$
@@
: Evaluation function could be matching-based or generative-based on the basis of different ways of utilizing LLMs. As shown in Figure 2, matching-based methods measure the semantic equivalence between the reference and hypothesis or measure the proper degree between the source text and hypothesis. Several works measure the semantic equivalence between the reference and hypothesis by using token-level matching functions in distributed representation space Zhang et al. (2020); Zhao et al. (2019) or discrete string space Lin (2004); Papineni et al. (2002). Others focus on sequence-level, such as Sellam et al. (2020); Rei et al. (2020). In contrast, generative-based methods include methods where LLMs are employed to generate evaluation metrics directly. These methods leverage the generative capabilities of LLMs to assess the quality of generated text by designing instructions.
Scope of this paper:
@@
Recent matching-based methods are based on a neural encoder to calculate a score-specific aspect of evaluation. However, these methods often face challenges such as limited interpretability, lower correlation with human judgments, and a restricted range of evaluated aspects Xu et al. (2023); Fu et al. (2023). Fortunately, the emerging capabilities of LLMs open up a wealth of possibilities for NLG evaluation. This includes improved interpretability through CoT, higher customization via instruction-following capabilities, and better alignment with human evaluations through RLHF Xu et al. (2023); Zheng et al. (2023). Given the abundance of recent surveys primarily focusing on matching-based evaluation methods (refer to Celikyilmaz et al. (2020); Sai et al. (2022); Goyal et al. (2023) for comprehensive summaries), our paper is dedicated to exploring more burgeoning generative-based methods. Figure 3 presents our taxonomy of generative-based evaluation. We classify relevant works into two main categories: prompt-based and tuning-based evaluation, depending on whether the LLM is tuned. Further, we divide these methods into subcategories: score-based, probability-based, likert-style, pairwise comparison, ensemble, and advanced evaluation protocols, each distinguished by their evaluation form. These categories will be detailed in Section 3.
@@ 3 Generative Evaluation
This section delves into a detailed exploration of these two overarching categories of evaluation methods, each accompanied by their respective evaluation protocols. Table 3.1 provides a comprehensive overview of current prompt-based and tuning-based evaluation methods. The table delineates their respective adaptation tasks, backbone models, scoring protocols, and evaluated aspects for clarity and reference.
@@ 3.1 Prompt-based Evaluation
Prompt-based text evaluation stands at the forefront of advancements in NLG, particularly leveraging the capabilities of LLMs. In this method, the evaluation process is intricately woven into the crafting of prompts – specialized cues designed to guide LLMs in assessing the quality and coherence of generated text. More recently, the Eval4NLP workshop held a shared task on prompting LLMs as explainable metrics Leiter et al. (2023). Typically, a prompt template serves as a structured framework that encompasses instructions, aspects, criteria, and desired output formats, providing a systematic guide for evaluating generated text. These templates empower researchers and practitioners to articulate precise evaluation requirements, ensuring consistency and reproducibility in the assessment process. By harnessing the prowess of LLMs, prompt-based evaluation not only provides a comprehensive understanding of NLG system performance but also offers a nuanced approach to extracting valuable insights.
@@
As the actual evaluation process often involves collaborative evaluation by multiple human annotators, some works utilize multiple LLM evaluators with different base models or prompts and allow them to evaluate the quality of generated text from different perspectives, as shown in Figure 5. Wu et al. (2023a) set multiple roles for the LLM to evaluate the quality of the generated summary by comparing it with the reference one on both subjective and objective dimensions. This work generates dynamic role profiles according to input texts and synthesizes the results of multiple roles as the final evaluation result. Li et al. (2023c) utilized multiple LLM evaluators to conduct a pairwise evaluation for the model-generated responses by performing multiple rounds of discussions on the comparison results to reach a mutual agreement on the pairwise scoring. Similarly, Zhang et al. (2023b) proposed to set up LLM evaluators as multiple-layer neural network structures. The bottom evaluators obtain the evaluation result of model-generated responses from a specific evaluation perspective. The upper evaluators receive all evaluation information from the previous layer and discuss it with each other to obtain a more comprehensive evaluation result. Besides, Chan et al. (2023) designed diverse communication strategies with various role prompts during collaborative discussions to evaluate pairwise generated responses.
@@
Some recent works investigate advanced evaluation techniques aimed at achieving more thorough and nuanced assessment outcomes by leveraging chain-of-thought, in-context learning capabilities, fine-grained analysis, etc. A representative fine-grained evaluation method is shown in Figure  4. Liu et al. (2023c) utilized LLMs with chain-of-thought (CoT) prompting and a form-filling paradigm to evaluate the quality of generated texts across various NLG tasks. Min et al. (2023) proposed a find-grained evaluation schema that first extracts a series of atomic facts from the LLM-generated long text, and then utilizes LLM evaluator to validate each atomic fact with the given knowledge source. Lu et al. (2023) proposed a new prompting method called Error Analysis Prompting (EAPrompt) that combines CoT to prompt the LLM evaluator to analyze different types of pre-defined errors (e.g., major and minor errors) in the generated translation based on the given source text and reference translation, and then measures the quality of a generated translation with the previous error analysis. To enhance and improve the robustness of LLM-based evaluators, Hasanbeig et al. (2023) proposed ALLURE, a systematic protocol for auditing and improving LLM-based evaluation with iterative in-context-learning. Considering that the evaluation with a single or few references may not accurately reflect the quality of the model’s hypotheses, Tang et al. (2023) leveraged LLMs to paraphrase a single reference into multiple high-quality ones in diverse expressions, which enhances various evaluation methods on MT, TS, and caption tasks. To further task advantages of the in-context learning capability of LLMs, Liu et al. (2023f) proposed AUTOCALIBRATE to automatically align and calibrate an LLM-based evaluator through incorporating the mined and calibrated rubrics into scoring instructions.
@@ 3.2 Tuning-based Evaluation
Recently, researchers increasingly turn their attention towards fine-tuning open-source language models (e.g., LLaMA), in lieu of traditional closed-based LLMs (e.g., ChatGPT and GPT-4). This shift is propelled by a thorough exploration of key perspectives, including the expenses associated with API calls, the robustness of prompting, and the pivotal consideration of domain adaptability.
@@
Typically, tuning-based methods construct evaluation data manually Zheng et al. (2023) or with the assistance of advanced LLMs (e.g., GPT-4) Xu et al. (2024), followed by performing supervised tuning. Similar to prompting-based evaluation, tuning-based methods can also be categorized into various types based on their scoring protocol, such as likert-style evaluation, probability-based evaluation and pairwise evaluation. In addition, based on the output explanations in supervised fine-tuning, these methods can be further divided into holistic evaluation or error-oriented evaluation. We will begin by introducing various types of scoring protocols and subsequently provide a summary of two output explanations in the final advance evaluation.
@@
Some works tune LLMs to provide quality ratings or labels for generated texts. Gekhman et al. (2023) employed FLAN-PaLM 540B Chung et al. (2022) to annotate the quality of real model-generated summaries and utilized these annotated data as training data to tune a light-weight LLM (e.g., T5-11B) as a factual consistency summary evaluator, which predicts “1" if the summary demonstrates factual consistency and “0" otherwise. Yue et al. (2023) reused and repurposed the existing fact-checking, NLI, and summarization tasks datasets and obtained simulated data from open-domain QA datasets to tune light-weight LLMs for attribution evaluation, which generates attributable, extrapolatory or contradictory labels for the generated answer with given query and reference documents. Li et al. (2023a) created a dataset containing multiple scenarios and used GPT-4 OpenAI (2023) to generate evaluation judgments for each scenario as supervision signals to tune LLaMA as a generative evaluator, which can output overall quality rating for individual LLM-generated response in various scenarios. Wang et al. (2023a) repurposed existing datasets with proper anonymization and new personalized labels to tune LLaMA2 Touvron et al. (2023) as a personalized story evaluation model which provides personalized evaluation for generated texts through outputting a grade in [1, 10] and detailed reviews. Kim et al. (2023a) prompted GPT-4 to construct training data, including reference answers and crafted diverse customized score rubrics, and used them to tune LLaMA to evaluate model-generated responses of given instruction, which is generalized to realistic user demands. Ke et al. (2023) instructed GPT-4 to collect referenced and reference-free training data with dialogue-based prompting, utilized to tune LLMs for evaluating the alignment of model-generated texts with human instructions through generating scores and explanations. Liu et al. (2023a) constructed a reference-free instruction-tuning dataset tailored for multi-aspect evaluation across summarization, dialogue and data-to-text tasks. Considering that there is an internal correlation between the evaluation aspects, this work tuned with auxiliary aspects additionally on diverse evaluation task forms. During inference, this work combined auxiliary and target aspects and predicted either the “Yes” or “No” label to judge whether the generated text satisfied the target aspect and compute the evaluation score.
@@
Some works train generative LLMs to calculate the generation probability of generated texts to evaluate text quality. For example, Thompson and Post (2020) trained a transformer as a multilingual reference-to-candidate paraphraser to obtain the generated probability of model-generated translation based on their reference texts. Qin et al. (2022) tuned the T5 model in the generative and discriminative fashion, and used the probability of generating a text as the quality score.